OpenStudy (anonymous):

What is the sum of an 8-term geometric series if the first term is −11, the last term is 180,224, and the common ratio is −4? @perl

OpenStudy (anonymous):

@kc_kennylau

OpenStudy (kc_kennylau):

Let \(a\) be the first term, \(r\) be the common ratio and \(n\) be the number of terms. The sequence is: \(\Large a, ar, ar^2, ar^3,\cdots,ar^{n-1}\) The series is: \(\Large \dfrac{a(r^n-1)}{r-1}\) Proof: http://fym.la.asu.edu/~tturner/MAT_117_online/SequenceAndSeries/geometric_sequence_17.gif

OpenStudy (kc_kennylau):

P.S. Series is the sum.

OpenStudy (anonymous):

isnt the formula \[a _{1-a _{n}r}/1-r\]

OpenStudy (anonymous):

@kc_kennylau

OpenStudy (kc_kennylau):

They're just the same :)

OpenStudy (kc_kennylau):

\[\Large\dfrac{a(r^n-1)}{r-1}=\dfrac{a(1-r^n)}{1-r}\]

OpenStudy (anonymous):

i got it wrong i get something out of this world idk why it isnt working

OpenStudy (kc_kennylau):

Mind showing me what you got? :)

OpenStudy (anonymous):

ok \[-11-(8)(-4)/1-4\]

OpenStudy (anonymous):

right this is how it looks when i pllugg in the numbers

OpenStudy (kc_kennylau):

Sorry I'm back now

OpenStudy (kc_kennylau):

Do you mind drawing it out? I cannot comprehend your LaTeX. Sorry for any inconvenience caused.

OpenStudy (anonymous):

:(can you just show me the proper wat and illl tell you my mistyake

OpenStudy (kc_kennylau):

OK :D \[\Large S=\dfrac{-11[(-4)^8-1]}{(-4)-1}\]

OpenStudy (anonymous):

why the exponent 8?

OpenStudy (kc_kennylau):

Number of terms :)

OpenStudy (anonymous):

and -1?

OpenStudy (anonymous):

ohh i see my mistake now i wasnt plugging it in properly ok now i will simplify

OpenStudy (kc_kennylau):

\[\Large a=-11,r=-4,n=8\] \[\Large S=\frac{a(r^n-1)}{r-1}=\frac{-11[(-4)^n-1]}{-4-1}\]

OpenStudy (kc_kennylau):

8* OK :)

OpenStudy (anonymous):

-144181?

OpenStudy (anonymous):

@kc_kennylau

OpenStudy (kc_kennylau):

lolz tagging me has no use xP

OpenStudy (anonymous):

lol you keep scaring me when you dont reply this website make you insecure

OpenStudy (anonymous):

is it right tho?

OpenStudy (kc_kennylau):

but I calculated 144177... http://bit.ly/1aCX0MU

OpenStudy (anonymous):

omggggggggggggggggggggggggggggggggggggggggggggggg

OpenStudy (anonymous):

thats was sooooo cool how did you do tht

OpenStudy (kc_kennylau):

http://lmgtfy.com

OpenStudy (anonymous):

http://lmgtfy.com/?q=2%2B2

OpenStudy (kc_kennylau):

gj :D

OpenStudy (anonymous):

lol